My problem is "9/14y = 3/7". I'm a bit confused with fractions and variables together. An explanation would be nice but I can probably attempt to figure it out if I have an answer.

Answer by erika514(39)

Okay this one is simple....
First we are going to divide by 9/14 or multiply by the reciprocal, 14/9 to both sides. By doing this, we should get y on one side (...because the fractions cancel) and on the other side, we need to multiply 3/7 times 14/9(our reciprocal. We can simplify these fractions to make them easier to multiply...we do this by simplifying numbers diagonal to each other. So we can simplify 3 and 9 by dividing by 3 for each number. 3 simplifies to 1(3/3=1) and 9 simplifies to 3 (9/3=3). We can also simplify the 7 and 14 by dividing by 7 for each number. 7 simplifies to 1(7/7=1) and 14 simplifies to 2(14/7=2).
Therefore our final answer is y=2/3. Hope this helps. Let me know if you have any questions. :)